Problem 3. Some bond calculus. (Hint: for this exercise, the formula for a geometric series might be useful: if q 1, then 1-q I will assume in the future, that this formula is well known. 1. Suppose a coupon bond has a face value F0 to be repaid after n years and promises to pay a coupon C equal to an interest rate of r percent of that face value for all years t = 1, . . . ,n (Note: I am using the notation x rather than i for the interest rate: that seems easier to read.). Currentlgy, ie in year t = 0, it is trading at price P

(a) Given n, find an equation, expressing the relationship between P/F on the left hand side and an expression in i and x/i on the right hand side, assuming 0 b) For what P does the yield equal the interest rate? c) Suppose n-oo and i >0. What is the formula in that case? Briefly describe, how i relates to P. For F-1000 and x = 5%, plot the yield i as a function of the price P, as the price varies from P 500 to P 1500.
Answer #1

Part (a)

Face Value = F

Coupon, C = x% of F = xF

Period from t=1 to t=n

Hence, Price of the bond today, P = Present value of all future coupons + Present value of face value at the time of maturity. Hence,

k=1

Hence

P T -n

--------------------

Part (b)

If yield = coupon rate, this means i = x

In that case,

P i -n

Hence, P = F

Hence, when current price is equal to face value of the bond, in that case yield = coupon rate.

-----------------

Part (c)

If n o infty then

(1+i)^{-n} o �

Hence, P / F = x / i

The relationship between P and i:

  1. Inverse: When i increases, P decreases and vice versa
  2. Non linear: The changes in P are disproportionate to changes in i

When F = 1000, x = 5% we can generate the following table

Price, P Yield, i = xF / P
500 10.00%
750 6.67%
1,000 5.00%
1,250 4.00%
1,500 3.33%

Please see the plot below: